Use inductive reasoning to predict the next number in each of the following list.




2, 5, 10, 17, 26, ?

1
Expert's answer
2022-03-17T03:46:38-0400

Solution

The logic behind the terms of the given series is simple, the first term is 2 and the second term is 5

"2 + 3 = 5"

The third term is 10

"5 + 5 = 10"

The fourth term is 17

"10 + 7 = 17"

The fifth term is 26

"17 + 9 = 26"


Hence, we are adding consecutive odd numbers starting from 3 to each preceding term.

The next number to be added to 26 will be 11

"26 + 11 = 37"


The next two consecutive odd numbers after 11 is 13 and 15 .Thus, the next 2 numbers in the series will be:

"37 + 13 = 50"

"50 + 15 = 65"


Thus, the next three terms in 2, 5, 10, 17, 26 ... are 37, 50, and 65.